OpenStudy (anonymous):

Are x = √6 and x=-√6 solutions to x²-3x+1 = 0? Explain your answer

OpenStudy (anonymous):

Using the quadratic formula \[x=[-b \pm \sqrt{b ^{2}-4ac} ]/2a\] where a = 1; b= -3 and c = 1 we get two answers; x=\[3/2 \pm \sqrt{5}/2\], so no

OpenStudy (anonymous):

if \[\sqrt{6} , -\sqrt{6}\] are roots then its equation is \[(x-\sqrt{6})(x+\sqrt{6})=0\] \[x^2-6=0\]